OpenStudy (anonymous):

how would you graph 10y-30=-6x? please an explaination of how to put this is in y intercept form

OpenStudy (anonymous):

10y-30=-6x 10y=-6x+30 y=(-6x+30)/10 y=(-3x+15)/5 y=-3x/5+3 that's in the y=mx+c form :)

OpenStudy (anonymous):

From the y intercept form that Zara showed you, you have your first point, which is the y intercept. y= -3x/5 + 3 +3 is the y intercept, which is where the line will cross on the Y plain Since we know the y intercept, which is (0,3) and we know our slope, which is -3x/5, we can start at the point (0,3) and then rise -3 and run 5 so our next point would be (5,0) Now that we have our two points (0,3) and (5,0) put a dot on the chart at these points and draw a line